# SlimLM: An Efficient Small Language Model for On-Device Document Assistance

### v1

@@ 1 Introduction
Our results show that SlimLM models perform comparably or even better than existing SLMs of similar sizes across standard metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), Semantic Textual Similarity (STS), Self-BLEU Zhu et al. (2018) for text diversity and GEval Liu et al. (2023). The smallest model SlimLM-125M demonstrates efficient performance on S24, making it suitable for widespread deployment. Larger variants, up to 1B parameters, offer enhanced capabilities while still operating within mobile constraints. To demonstrate real-world applicability, we develop an Android application showcasing SlimLM’s document assistance capabilities [^1] Code and data will be released at anonymous.slimlm (Sec. 4).
@@ 2.2.2 Data Annotation
We propose a novel approach for annotating documents using GPT-4o-mini OpenAI (2024) to generate comprehensive annotations for three key tasks in DocAssist: SUMM, QS, and QA. For each document, our method produces five distinct examples: one summary, one set of three suggested questions, and three question-answer pairs.
@@
Our approach serves several crucial functions: it facilitates intent classification training, enables task-specific response generation, enhances contextual understanding, ensures versatility in document handling, and maintains quality control in annotations. By leveraging the capabilities of GPT-4o-mini, we aim to generate high-quality annotations that capture the nuances and complexities of the documents. The in-context examples and detailed requirements are provided in Tables 12, 9, 10 and 11.
@@
Table 4 provides insight into the token usage statistics for the GPT-4o-mini model in annotating 82,850 documents. The relatively low standard deviation in completion tokens suggests consistent-length responses across different documents, which is desirable for maintaining annotation quality and consistency. The annotation process yields $\sim$414K examples for DocAssist. Of these, $\sim$2K examples were randomly selected for the test set, with the remaining examples allocated to the training set.
@@
Table 4: Token usage statistics for GPT-4o-mini model in annotating 82,850 documents.
@@ 2.3.2 Fine-tuning
Following pre-training, we fine-tuned the models on the training set of DocAssist that comprises $\sim$412K examples to enhance document assistance capabilities by teaching them to handle specific tasks based on user requests. The process instructs the model to first identify the appropriate task from the user’s input and then generate a response that match the quality of GPT-4o-mini for the identified task. The fine-tuning loss function is also an autoregressive objective, defined as:
@@
where $x$ is the input sequence (system prompt, document and user request), $y_{i}$ is the $i^{th}$ token in the target response generated by GPT-4o-mini, $y_{<i}$ denotes all tokens preceding $y_{i}$ in the target response $m$ is the length of the target response.
@@ 3.2 Results
As model size increases (Table 5), we observe consistent improvement across all metrics, suggesting good scalability. Our largest model, SlimLM-1B, approaches the performance of the much larger model Qwen2-1.5B-Instruct, highlighting the potential for SlimLM to achieve competitive results with reduced computational requirements. While GPT-4o-mini still leads in overall performance, our SlimLM models offer a range of efficient options for various computational constraints and privacy concerns in document assistance tasks.
@@ 4 Use Case
Adobe’s Acrobat relies on API calls for SUMM, QS and QA, incurring significant costs and raising privacy concerns. SlimLM can be deployed into the Acrobat mobile apps, enabling local processing of documents. This approach eliminates the need for external API calls, substantially reducing operational costs while enhancing user privacy by keeping document content on the device.
@@ Appendix A Appendix
Table 9: {{summ_req}}. Instructional prompt designed to guide GPT-4o-mini how to summarize the document contents.
@@
Table 10: {{qa_req}}. Instructional prompt designed to guide GPT-4o-mini how to answer questions for the Q/A task.
@@
Table 11: {{suggestion_req}}. Instructional prompt designed to guide GPT-4o-mini how to generate suggested questions for a given document. The suggested questions aims to guide users what should be asked to understand the document.
@@
Examples: DOCUMENT CONTEXT (may be truncated) XYZ Corporation has released its Q3 earnings report. The company reported a net income of \$5 million, up from \$3 million in Q2. The earnings per share (EPS) increased to \$0.25 from \$0.15 in the previous quarter. The revenue for Q3 was \$50 million, a 10% increase compared to Q2. The company’s operating expenses increased by 5% to \$20 million. The table below summarizes the key financial figures: Metric Q3 Q2 % Change Net Income \$5M \$3M 66.67% EPS \$0.25 \$0.15 66.67% Revenue \$50M \$45M 11.11% Operating Expenses \$20M \$19M 5.26% RESPONSE
@@
    "summarization": "XYZ Corporation’s Q3 earnings report shows significant growth with net income rising to \$5 million from \$3 million in Q2.
    EPS increased to \$0.25, revenue grew by 11.11% to \$50 million, and operating expenses rose by 5.26% to \$20 million.",
@@
      "What’s the Q2 to Q3 revenue percentage increase?",
      "What’s the EPS increase from Q2 to Q3?",
      "What’s the operating expenses percentage change between Q2 and Q3?"
@@
      "The Q2 to Q3 revenue percentage increase is 11.11%.",
      "The EPS increased from \$0.15 in Q2 to \$0.25 in Q3, which is a \$0.10 or 66.67% increase.",
      "The operating expenses percentage change between Q2 and Q3 is 5.26%."
DOCUMENT CONTEXT (may be truncated) # Document Excerpt 1: In the land of X, fairies and trolls have always been at odds. The fairies, known for their beauty and grace, live in the enchanted forest, while the trolls, known for their strength and cunning, reside in the dark caves. The two groups have had friction for centuries, with each side trying to outwit… # Document Excerpt 2: The fairies use their magical powers to protect their forest, while the trolls rely on their dancing abilities to enter the fairies’ territory. The story of their ongoing struggle is a tale of cunning, bravery, and the power of friendship. RESPONSE
@@
    "summarization": "In the land of X, fairies and trolls have been in conflict for centuries. Fairies, living in the enchanted forest, use magic
    for protection, while trolls from dark caves use dancing to infiltrate fairy territory. Their struggle is characterized by cunning, bravery,
    and unexpected friendship.",
@@
      "What traits define fairies and trolls?",
      "How do fairies defend their forest?",
      "What strategies do trolls use for entering the fairies’ territory?"
@@
      "Fairies are defined by their beauty and grace, while trolls are known for their strength and cunning.",
      "Fairies use their magical powers to protect their forest.",
      "Trolls rely on their dancing abilities to enter the fairies’ territory."

### v2

@@ 1 Introduction
Our results show that SlimLM models perform comparably or even better than existing SLMs of similar sizes across standard metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), Semantic Textual Similarity (STS), Self-BLEU Zhu et al. (2018) for text diversity and GEval Liu et al. (2023). The smallest model SlimLM-125M demonstrates efficient performance on S24, making it suitable for widespread deployment. Larger variants, up to 1B parameters, offer enhanced capabilities while still operating within mobile constraints. To demonstrate real-world applicability, we develop an Android application showcasing SlimLM’s document assistance capabilities (Sec. 4).
@@ 2.2.2 Data Annotation
We propose an approach for annotating documents using a commercial LLM to generate comprehensive annotations for three key tasks in DocAssist: SUMM, QS, and QA. For each document, our method produces five distinct examples: one summary, one set of three suggested questions, and three question-answer pairs.
@@
Our approach serves several crucial functions: it facilitates intent classification training, enables task-specific response generation, enhances contextual understanding, ensures versatility in document handling, and maintains quality control in annotations. By leveraging the capabilities of the commercial LLM, we aim to generate high-quality annotations that capture the nuances and complexities of the documents. The in-context examples and detailed requirements are provided in Tables 12, 9, 10 and 11.
@@
Table 4 provides insight into the token usage statistics for the commercial LLM in annotating 82,850 documents. The relatively low standard deviation in completion tokens suggests consistent-length responses across different documents, which is desirable for maintaining annotation quality and consistency. The annotation process yields $\sim$414K examples for DocAssist. Of these, $\sim$2K examples were randomly selected for the test set, with the remaining examples allocated to the training set.
@@
Table 4: Token usage statistics for the commercial LLM in annotating 82,850 documents.
@@ 2.3.2 Fine-tuning
Following pre-training, we fine-tuned the models on the training set of DocAssist that comprises $\sim$412K examples to enhance document assistance capabilities by teaching them to handle specific tasks based on user requests. The process instructs the model to first identify the appropriate task from the user’s input and then generate a response that match the quality of the commercial LLM for the identified task. The fine-tuning loss function is also an autoregressive objective, defined as:
@@
where $x$ is the input sequence (system prompt, document and user request), $y_{i}$ is the $i^{th}$ token in the target response generated by the commercial LLM, $y_{<i}$ denotes all tokens preceding $y_{i}$ in the target response $m$ is the length of the target response.
@@ 3.2 Results
As model size increases (Table 5), we observe consistent improvement across all metrics, suggesting good scalability. Our largest model, SlimLM-1B, approaches the performance of the much larger model Qwen2-1.5B-Instruct, highlighting the potential for SlimLM to achieve competitive results with reduced computational requirements. While the commercial LLM still leads in overall performance, our SlimLM models offer a range of efficient options for various computational constraints and privacy concerns in document assistance tasks.
@@ 4 Use Case
SlimLM can be deployed into the mobile apps, enabling local processing of documents. This approach eliminates the need for external API calls, substantially reducing operational costs while enhancing user privacy by keeping document content on the device.
@@ Appendix A Appendix
Table 9: {{summ_req}}. Instructional prompt designed to guide the commercial LLM how to summarize the document contents.
@@
Table 10: {{qa_req}}. Instructional prompt designed to guide the commercial LLM how to answer questions for the Q/A task.
@@
Table 11: {{suggestion_req}}. Instructional prompt designed to guide the commercial LLM how to generate suggested questions for a given document. The suggested questions aims to guide users what should be asked to understand the document.
@@
Examples: DOCUMENT CONTEXT (may be truncated) ### Document Excerpt 1: English Language Arts, Grade 9 Inquiry question: How does an individual address injustice in society? Abstract This sample highlights my thinking and planning as it relates to a poetry lesson based on the redesigned BC Curriculum. I have made explicit connections to the First Peoples Principles of Learning, the Big Ideas, the Curricular Competencies and Content learning standards, Core Competencies, and. I have deliberately chosen activities that offer opportunities for multiple entry points, choice, and differentiation. I have also included some samples of student work. Through the exploration of text, discussion, and the activities of the lesson, students have opportunities to respond to the inquiry question with an informed and thoughtful voice. As students engage with text, collaborate with one another, create personal responses to text, and finally reflect on the process of their learning, they are doing so purposefully, considering the inquiry question. Student learning depends on considering prior knowledge, sharing and developing ideas, and revising and extending thinking. A(truncated due to length…) ### Document Excerpt 2: Understand text differently depending on their worldviews and perspectives. Texts are socially, culturally, and historically constructed. Questioning what we hear, read, and view contributes to our ability to be educated and engaged citizens. Curricular Competencies. Think critically, creatively, and reflectively to explore ideas within, between, and beyond texts. Construct meaningful personal connections between self, text, and world. Respond to text in personal, creative, and critical ways. Use writing and design processes to plan, develop, and create engaging and meaningful literary and informational texts for a variety of purposes and audiences Content Strategies and processes: o Writing processes Core Competencies The Learning Experience Exploring the concept of injustice Students are invited to bring to class examples of injustice in our community and/or in our world. This provides multiple entry points into the inquiry for the students. These examples provide opportunities for students to discuss the diverse ways in which injustice is evident/experienced in their(truncated due to length…) RESPONSE
@@
    "summarization": "The document outlines a Grade 9 poetry lesson focused on addressing injustice, using inquiry-based activities aligned with
    the redesigned BC Curriculum and emphasizing critical thinking, personal connections, and creative expression.",
@@
        "What inquiry question is central to the lesson?",
        "How do students engage with the concept of injustice?",
        "What is the purpose of the empathy journal?"
@@
        "The central inquiry question is: How does an individual address injustice in society?",
        "Students engage with the concept of injustice by discussing real-world examples and exploring texts that highlight these issues.",
        "The purpose of the empathy journal is to help students transform negative messages into positive ones and reflect on the emotional impact of
        those transformations."
DOCUMENT CONTEXT (may be truncated) ### Document Excerpt 1: Demographic situation in Bulgaria in 2018: Number of population decreases and population ageing continue; Unbalanced territorial distribution of population goes deeper; Number of live births and crude birth rate decrease; Number of deaths and crude death rate decrease; Infant mortality decreases; Number of marriages and number of divorces increase; Life expectancy at birth is unchanged. Population number and structures 1 As of 31 December 2018 the population of Bulgaria is 7 000 039 persons representing 1.4% of the EU population. Compared to 2017, the country population decreases by 49 995 persons or by 0.7%. Male population was 3 395 701 (48.5%) and female - 3 604 338 (51.5%) or to 1 000 males correspond 1 061 females. Males prevail among the population aged up to 53(truncated due to length…) ### Document Excerpt 2: the end of 2018, the number of population over working age is 1 732 thousand persons or 24.8% and under working age are 1 067 thousand persons or 15.2% of the country population. 1. Population under, at and over working age as of 31.12. Year Age Under working age Working age Over working age (Per cent) 1990 21.6 55.5 22.9 1995 19.1 56.6 24.3 2001 16.3 59.2 24.5 2005 14.8 62.4 22.8 2010 14.6 62.7 22.7 2011 14.3 61.7 24.0 2012 14.4 61.8 23.8 2013 14.6 (truncated due to length…) RESPONSE
@@
    "summarization": "In 2018, Bulgaria’s population declined to 7 million, with notable demographic changes including decreased birth rates,
    increased divorces, and a high mortality rate compared to EU averages.",
@@
        "What was Bulgaria’s population at the end of 2018?",
        "How did the birth and death rates change?",
        "Which municipality had the highest population in 2018?"
@@
        "At the end of 2018, Bulgaria’s population was 7,000,039.",
        "Both the number of live births and deaths decreased in 2018.",
        "The municipality with the highest population in 2018 was Stolichna, with 1,328,120 residents."